Solve u, = 4 for 0 5xs1, given u(0,t) = 0, u,(1,t) = 0, u(x,0)=1. 2 fomu sin 1 Answer: u(x,t) = { e sin( n +) 7x 0 1 +
Answer #1

Ans) 21+ = Uxx (0,t)=0 2011, t= 0 2 x (1,4)=0 21(0,0)=1 Let Ulut) = XENITIEI Ulot = X (0) T (t) = 0 = X(0)=0 Ux(x,x) = x (x 12 X (1)= 0 =) Costo 1 = 0 5- di = 1n + 2 ) for n=0, 1,2- 2 2 n=0,), 2 - - T - d = (n+1) Xncns - Bor Sin Int / TX -(nt 21 17²On = 2 S 1. In[n+4) mx] de 3 - - 2 Cos (nt I) at 1h+sho =-9 (os/2+1/2 ). Tl - Cosoo Int - 2 O ht 50 23 tit 21/n, t 9 Sin [Int